OpenStudy (anonymous):

If f(2)=6, f'(2)=-4, and f"(2)=2, what is (d^2*(f^3 (x)))/(dx^2 ) at x=2?

myininaya (myininaya):

\[\frac{d^2([f(x)]^3)}{dx^2}\] \[=\frac{d}{dx}(3[f(x)]^2f'(x))=3[2f(x)f'(x)f'(x)+[f(x)]^2f''(x)]\]

myininaya (myininaya):

\[3[f(2)f'(2)f'(2)+[f(2)]^2f''(2)]\]

myininaya (myininaya):

now plug in and evaluate and simplify

OpenStudy (anonymous):

Thank you I'm going to try it now.

myininaya (myininaya):

I used chain rule and product rule by the way

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

The answers I keep coming up don't seem correct. Can you please explain in more detail.

myininaya (myininaya):

lol and I didn't notice you were having problems with this money sorry

OpenStudy (amistre64):

just like a woman, they state their opinions and then yougotta live with em ;)

myininaya (myininaya):

\[3[6(-4)(-4)+(6)^2(2)]=3[6(16)+36(2)]=3[96+72]=3[168]\]

myininaya (myininaya):

= 504 is what I'm getting

OpenStudy (amistre64):

ipso facto

OpenStudy (anonymous):

I was missing the 2 in front of the f(x). But I understand now.

OpenStudy (amistre64):

i dont think id of come to that today;

myininaya (myininaya):

lol poor amistre do you ever come to it? ;)

OpenStudy (anonymous):

The radius of a sphere is increasing at a constant rate of 2cm/sec. At the instant when the volume of the sphere is 36π cm^3, what is the rate that the surface area is increasing? Surface area=4π r^2 Volume=4/3π r^3 This one I need a full full explanation on

myininaya (myininaya):

\[V(t)=\frac{4}{3} \pi [r(t)]^3=> V'(t)=\frac{4}{3} \pi 3 [r(t)]^2 r'(t)\]

myininaya (myininaya):

\[\S(t)=4 \pi [r(t)]^2 => \S'(t)=4 \pi 2r(t) r'(t)=8 \pi r(t)r'(t)\]

myininaya (myininaya):

\[V'(t)=4 [r(t)]^2 r'(t)\]

myininaya (myininaya):

so r'=2 do you see that in the first sentence?

myininaya (myininaya):

\[V=36 \pi\] from second sentence

myininaya (myininaya):

we want to know S' based on second sentence

myininaya (myininaya):

\[36=\frac{4}{3} \pi r^3\] solve this for r

myininaya (myininaya):

to find S' this is the only thing we need since r' is already given

myininaya (myininaya):

oops the V=36 pi

myininaya (myininaya):

\[36 \pi =\frac{4}{3} \pi r^3\] *

myininaya (myininaya):

solve that for r

OpenStudy (anonymous):

So am I suppose to divide both sides by 4/3\[\pi\] ?

myininaya (myininaya):

or multiply both sides by 3/(4pi)

OpenStudy (anonymous):

so would I have 27pi=r^3?

myininaya (myininaya):

well the pi's would cancel

OpenStudy (anonymous):

27=r^3?

myininaya (myininaya):

yes since pi/pi=1 so r=3

myininaya (myininaya):

\[\S'=8 \pi r r'\]

myininaya (myininaya):

and remember r'=2

myininaya (myininaya):

So now you can find S'

OpenStudy (anonymous):

So I should come up with 48pi cm^2/sec correct?

myininaya (myininaya):

48 pi is right!

OpenStudy (anonymous):

Thanks!!